A large 3-bladed fan has a radius of 3.5 m. Each blade has a mass of 130 kg. Treat each blade as a thin rod. Determine the angular momentum of the fan in unit vector notation if it spins at 350 rpm. Assume +x is to the right and +y is up.
Answer #1

here,

the length of each blade , r = 3.5 m

mass of each blade , m = 130 kg

angular speed , w = 350 rpm = 36.62 rad/s

the angular momentum of the fan , L = (3 * moment of inertia of each blade * alpha

L = (3 * m * l^2 / 3 ) * w

L = 130 * 3.5^2 * 36.62 kg.m^2 /s

L = 5.83 * 10^4 kg.m^2 /s